OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} (1-\sqrt{x+1})/(x ^{2}+x)\]

OpenStudy (anonymous):

Not sure what to do with this

OpenStudy (anonymous):

Sorry to whoever answered this before my pc died

OpenStudy (anonymous):

usual gimmick is to rationalize the numerator

OpenStudy (anonymous):

you know what i mean by that? multiply top and bottom by \[1+\sqrt{x+1}\] then cancel

OpenStudy (anonymous):

leave the denominator in factored form, don't multiply it out you should get \[\frac{1-(x+1)}{(x^2+x)(1+\sqrt{x+1})}\] as a first step

OpenStudy (anonymous):

okay, ya thats what I had, but couldn't get past that.

OpenStudy (anonymous):

then numerator is \(-x\) which cancels since you can factor the denominator as \[x(x+1)(1+\sqrt{x+1})\]

OpenStudy (anonymous):

leaves you with \[\frac{-1}{(x+1)(1+\sqrt{x+1})}\] then replace \(x\) by \(0\) since you will not have a zero in the denominator anymore

OpenStudy (imstuck):

EXCELLENT JOB, @satellite73!!!

OpenStudy (anonymous):

Yes @satellite73 thank you very much!